# Theoretically Achieving Continuous Representation of Oriented Bounding Boxes

### v1

@@ 8.5 Details of Computing IoU scores
$r_{sx}$ $=\frac{1-\sqrt{1-4\cdot\frac{h^{2}}{w^{2}}\cdot r_{s}(1-r_{s})}}{2},r_{sy}=r_{s}$ (20)
$l_{1}$ $=\sqrt{(r_{sx}w)^{2}+(r_{sy}h)^{2}}$
$l_{2}$ $=\sqrt{[(1-r_{sx})w]^{2}+[(1-r_{sy})h]^{2}}$
$l_{3}$ $=\sqrt{(r_{sx}w)^{2}+[(1-r_{sy})h]^{2}}$
$l_{4}$ $=\sqrt{[(1-r_{sx})w]^{2}+(r_{sy}h)^{2}}$
@@
$\begin{split}I_{0,1}&=[1-\frac{(1-2r_{sx})r_{sx}w^{2}}{(1-r_{sy})h^{2}}]l_{1}l_{2}\\ IoU_{0,1}&=\frac{I_{0,1}}{l_{1}l_{2}+l_{3}l_{4}-I_{0,1}}\end{split}$ (21)
@@
$\begin{split}I_{0,2}&=(1-\frac{(1-2r_{sy})r_{sy}h^{2}}{(1-r_{sx})w^{2}})l_{1}l_{2}\\ IoU_{0,2}&=\frac{I_{0,2}}{l_{1}l_{2}+l_{3}l_{4}-I_{0,2}}\end{split}$ (22)
@@
$\begin{split}I_{0,3}&=\frac{(r_{sx}+r_{sy}-2r_{sx}r_{sy})^{2}}{(1-r_{sx})(1-r_{sy})}\times\frac{wh}{2}\\ IoU_{0,3}&=\begin{cases}\frac{I_{0,3}}{2l_{1}l_{2}-I_{0,3}}&I_{0,3}\neq 0\\ 0&I_{0,3}=0\end{cases}\end{split}$ (23)
@@
$\begin{split}h_{1}&=\frac{1}{2}w-\frac{\frac{1}{2}-r_{sy}}{1-r_{sy}}r_{sx}w\\ h_{2}&=\frac{1}{2}h-\frac{\frac{1}{2}-r_{sx}}{1-r_{sx}}r_{sy}h\\ \tan\alpha&=\frac{\frac{\frac{1}{2}-r_{sx}}{1-r_{sx}}l_{4}}{\frac{1}{2(1-r_{sy})}l_{3}}\quad\tan\beta=\frac{\frac{\frac{1}{2}-r_{sy}}{1-r_{sy}}l_{3}}{\frac{1}{2(1-r_{sx})}l_{4}}\\ I_{1,2}&=2\frac{\tan\alpha\tan\beta}{\tan\alpha+\tan\beta}(h_{1}^{2}+h_{2}^{2})+2h_{1}h_{2}\\ IoU_{1,2}&=\begin{cases}\frac{I_{1,2}}{2l_{3}l_{4}-I_{1,2}}&\tan\alpha\tan\beta\neq 0\\ \frac{2h_{1}h_{2}}{2l_{3}l_{4}-2h_{1}h_{2}}&\tan\alpha\tan\beta=0\end{cases}\end{split}$ (24)
@@
$M(w,h,r_{s})=\begin{bmatrix}1&IoU_{0,1}&IoU_{0,2}&IoU_{0,3}\\ IoU_{0,1}&1&IoU_{1,2}&IoU_{0,2}\\ IoU_{0,2}&IoU_{1,2}&1&IoU_{0,1}\\ IoU_{0,3}&IoU_{0,2}&IoU_{0,1}&1\\ \end{bmatrix}$ (25)
@@ 8.6 Proof of Encoding Continuity of COBB
To establish the continuity of IoU scores, we consider an OBB perturbed from $OBB$, denoted as $OBB^{\prime}$. The similarity between $x_{c}$, $y_{c}$, $w$, $h$, and $r_{s}$ of $OBB$ and $OBB^{\prime}$ emerges from their continuous evolution during shape transformations. Eq. 17 confirms the analogous construction of four OBBs using the parameters of $OBB$ and $OBB^{\prime}$. IoU scores represent the overlaps between $OBB$ and the four OBBs sharing the same $x_{c}$, $y_{c}$, $w$, $h$, and $r_{s}$ as $OBB$. As both $OBB$ and the associated four OBBs remain analogous before and after perturbation, minor disturbances on $OBB$ do not significantly alter its IoU scores.
@@ 8.8 Comparison of Different Regression Targets
$\Delta OBB=(1-IoU[f_{Dec}(r),f_{Dec}(r+\epsilon)])\epsilon^{-1}$ (26)
@@ 9.3 Futher Ablation Study for COBB
Models $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$
Oriented R-CNN [30] 78.73 55.07 50.57
+COBB-sig 79.09 55.61 50.80
+COBB-ln 79.23 56.15 50.55
$\lambda$ $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$
1 73.41 44.21 43.10
2 74.00 44.03 43.29
4 73.63 43.85 42.83

### v2

@@ 9.5 Details of Computing IoU scores
$r_{sx}$ $=\frac{1-\sqrt{1-4\cdot\frac{h^{2}}{w^{2}}\cdot r_{s}(1-r_{s})}}{2},r_{sy}=r_{s},$ (20)
$l_{1}$ $=\sqrt{(r_{sx}w)^{2}+(r_{sy}h)^{2}},$
$l_{2}$ $=\sqrt{[(1-r_{sx})w]^{2}+[(1-r_{sy})h]^{2}},$
$l_{3}$ $=\sqrt{(r_{sx}w)^{2}+[(1-r_{sy})h]^{2}},$
$l_{4}$ $=\sqrt{[(1-r_{sx})w]^{2}+(r_{sy}h)^{2}},$
@@
$\begin{split}I_{0,1}&=[1-\frac{(1-2r_{sx})r_{sx}w^{2}}{(1-r_{sy})h^{2}}]l_{1}l_{2},\\ IoU_{0,1}&=\frac{I_{0,1}}{l_{1}l_{2}+l_{3}l_{4}-I_{0,1}}.\end{split}$ (21)
@@
$\begin{split}I_{0,2}&=(1-\frac{(1-2r_{sy})r_{sy}h^{2}}{(1-r_{sx})w^{2}})l_{1}l_{2},\\ IoU_{0,2}&=\frac{I_{0,2}}{l_{1}l_{2}+l_{3}l_{4}-I_{0,2}}.\end{split}$ (22)
@@
$\begin{split}I_{0,3}&=\frac{(r_{sx}+r_{sy}-2r_{sx}r_{sy})^{2}}{(1-r_{sx})(1-r_{sy})}\times\frac{wh}{2},\\ IoU_{0,3}&=\begin{cases}\frac{I_{0,3}}{2l_{1}l_{2}-I_{0,3}}&I_{0,3}\neq 0,\\ 0&I_{0,3}=0.\end{cases}\end{split}$ (23)
@@
$\begin{split}h_{1}&=\frac{1}{2}w-\frac{\frac{1}{2}-r_{sy}}{1-r_{sy}}r_{sx}w,\\ h_{2}&=\frac{1}{2}h-\frac{\frac{1}{2}-r_{sx}}{1-r_{sx}}r_{sy}h,\\ \tan\alpha&=\frac{\frac{\frac{1}{2}-r_{sx}}{1-r_{sx}}l_{4}}{\frac{1}{2(1-r_{sy})}l_{3}},\quad\tan\beta=\frac{\frac{\frac{1}{2}-r_{sy}}{1-r_{sy}}l_{3}}{\frac{1}{2(1-r_{sx})}l_{4}},\\ I_{1,2}&=2\frac{\tan\alpha\tan\beta}{\tan\alpha+\tan\beta}(h_{1}^{2}+h_{2}^{2})+2h_{1}h_{2},\\ IoU_{1,2}&=\begin{cases}\frac{I_{1,2}}{2l_{3}l_{4}-I_{1,2}}&\tan\alpha\tan\beta\neq 0,\\ \frac{2h_{1}h_{2}}{2l_{3}l_{4}-2h_{1}h_{2}}&\tan\alpha\tan\beta=0.\end{cases}\end{split}$ (24)
@@
$M(w,h,r_{s})=\begin{bmatrix}1&IoU_{0,1}&IoU_{0,2}&IoU_{0,3}\\ IoU_{0,1}&1&IoU_{1,2}&IoU_{0,2}\\ IoU_{0,2}&IoU_{1,2}&1&IoU_{0,1}\\ IoU_{0,3}&IoU_{0,2}&IoU_{0,1}&1\\ \end{bmatrix}.$ (25)
@@ 9.6 Proof of Encoding Continuity of COBB
To establish the continuity of IoU scores, we consider an OBB perturbed from OBB $X$, denoted as $Y$. The similarity between $x_{c}$, $y_{c}$, $w$, $h$, and $r_{s}$ of $X$ and $Y$ emerges from their continuous evolution during shape transformations. Eq. 17 confirms the analogous construction of four OBBs using the parameters of $X$ and $Y$. IoU scores represent the overlaps between $X$ and the four OBBs sharing the same $x_{c}$, $y_{c}$, $w$, $h$, and $r_{s}$ as $X$. As both $X$ and the associated four OBBs remain analogous before and after perturbation, minor disturbances on $X$ do not significantly alter its IoU scores.
@@ 9.8 Comparison of Different Regression Targets
Table 8: NAE of parameters. $\alpha$ is the parameter for orientation determination in Gliding Vertex. In our method and Gliding Vertex xy denotes the center, wh represents width and height, $r_{s}$ and IoU scores are parameters of our method, and $\alpha$ denotes the additional parameters for orientation determination of Gliding Vertex.
Method xy wh $r_{s}$ IoU scores $\alpha$
Ours. 0.18% 0.31% 0.28% 0.41% -
Glidng Vertex 0.17% 0.31% - - 0.75%
@@
Models $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$
Oriented R-CNN [33] 78.73 55.07 50.57
+COBB-sig 79.09 55.61 50.80
+COBB-ln 79.23 56.15 50.55
$\lambda$ $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$
1 73.41 44.21 43.10
2 74.00 44.03 43.29
4 73.63 43.85 42.83
Table 12: Experiments on SOTA baselines on DOTA-v1.0.
Method $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$
SES 75.72 48.86 46.19
SES + Ours. 76.43(+0.71) 49.28(+0.42) 46.59(+0.40)
LSKNet-t 76.68 49.28 46.15
LSKNet-t + Ours. 77.29(+0.61) 50.91(+1.63) 47.62(+1.47)
$\Delta OBB=(1-IoU[f_{Dec}(r),f_{Dec}(r+\epsilon)])\epsilon^{-1}.$ (26)
9.9 Models’ Ability to Well Estimate Parameters
We contend that the parameters in our COBB are easily estimable due to their continuity. To elucidate this assertion, we introduce the Normalized Average Error (NAE) as a metric for assessing the difficulty of parameter estimation. Given the $i$-th prediction of the parameter as $P_{i}$ and its corresponding ground truth as $T_{i}$, the NAE is defined as $\textrm{NAE}=\frac{1}{N}\sum_{i=1}^{N}\frac{(P_{i}-T_{i})^{2}}{(\max(T)-\min(T))^{2}}$, where $N$ represents the number of predictions, and $\max(T)$ and $\min(T)$ denote the maximum and minimum values of ground truth values.
We posit that parameters with small NAE values are more ready to estimate. The NAE values of our method and Gliding Vertex are documented in Tab. 8. Without significantly influencing the prediction difficulty of the other parameters, $r_{s}$ and IoU scores in our method can be better estimated than $\alpha$s in Gliding Vertex.
@@ 10.3 Futher Ablation Study for COBB
10.4 Experiments on Latest Techniques
To verify the advantage of our method on SOTA methods, we added our method to the latest SOTAs, including SES [36] and LSKNet [12]. Experiments in Tab. 12 demonstrate that our method achieves a relatively large $\text{mAP}_{50}$ improvement, 0.71% and 0.61%, over SES and LSKNet, respectively.